OpenStudy (anonymous):

solve & graph the polynomial function. g(x)=x^4-4x^3+3x^2

OpenStudy (anonymous):

g(x) = (x^2)(x^2 - 4x + 3) = (x^2)(x - 3)(x - 1) So, zero at x = 0, 1, and 3 Graph attached.

OpenStudy (anonymous):

k im taking a look at it now

OpenStudy (anonymous):

thanks